OpenStudy (anonymous):

Solve for x: quantity x minus 3 all over 4 equals 2

OpenStudy (anonymous):



OpenStudy (anonymous):

x=11

OpenStudy (callisto):

@adityanaik28 Please show your workings instead of just giving answer. This act violates the Code of Conduct of this site. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

ya sure....... first x-3=2*4=8 thn x=8+3=11 get it??

OpenStudy (anonymous):

yess thankks